ereise 19abcd Qui Check 63 66 69 7275 78 81 8 87 90 93 Height (inches) Guided Part 1 How many classes are there? There are 11 What is the class width? The class width is 3 Part 2 Which class has the lowest frequency?. The class 93 96 has the lowest frequency. Show E classes. Link to T Tabl Formu Save for it Assi Part 3 out of 3 tage of students are less than 78 inches talh The percentage of students less than 78 inches tall is
HW 2.2 Frequency Distribution and Their Graphs Previous 123 4 Next Question 2 of 4 (1 point) View problem in a pop-up 2.2 Section Exercise 19ab 30 18 63 66 69 72 75 78 81 84 87 90 939 9 Height (inches) Part 1 How many classes are there? There are 11 classes.
Answer #1

The frequency distribution table for the given histogram will be as follows

Height Frequency
63 -66 7
66 - 69 6
69 - 72 5
72 - 75 12
75 - 78 16
78 - 81 17
81 - 84 13
84 - 87 10
87 - 90 9
90 - 93 5
93 - 96 1
Total 101

From the table we can find out that class width = Upper limit - Lower limit of any class = 69-66 = 3.

part - 3.

the students less than 78 inches tall = 7+6+5+12+16 = 46

percentage of students less than 78 inches tall = (46/101)*100 = 45.54%